OpenStudy (anonymous):

determine whether the graphs of the two equations are perpendicular y=9x-3 2y=3-x

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

Im sooooooooo sorry!

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

& they arent.

OpenStudy (saifoo.khan):

their product must be equal to -1

OpenStudy (anonymous):

sorry for what ?

OpenStudy (saifoo.khan):

for replying so late. i forgot that u had a ques. lol.

OpenStudy (anonymous):

can you still help me with another one

OpenStudy (saifoo.khan):

Sure!

OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point. Express your answer in the form x=a,y,=b or y =mx + b m=-4(7,0) the equation of the line is y =

OpenStudy (saifoo.khan):

y = mx + c 0 = -4(7) + c c = 28 y = mx + c y = -4x + 28

OpenStudy (anonymous):

lol can you be my tutor your good :)

OpenStudy (anonymous):

k I just got a few more if its ok

OpenStudy (saifoo.khan):

LoL, Ok, sure. I can b.. ;)

OpenStudy (anonymous):

k this one is hard and if you could explain it to me find the indicated outputs for f(x) =3x2-4x and the 2 by the x is supose to be a small two i think it means to the second power right ?

OpenStudy (saifoo.khan):

Infact, i do teach many ppl from Openstudy personally.

OpenStudy (saifoo.khan):

does "and the 2" means: 3x^2 -4x + 2 ??

OpenStudy (saifoo.khan):

Quadratic Eq.